OpenStudy (anonymous):

3 square root of 108 -2 square root of 18 -3 square root of 48

OpenStudy (anonymous):

hii there

OpenStudy (anonymous):

i need help on a math question

OpenStudy (anonymous):

hi :)

OpenStudy (anonymous):

so what is it that you want to ask?

OpenStudy (anonymous):

is that 3 * the square root of 108??

OpenStudy (anonymous):

here ill write the equation down

OpenStudy (anonymous):

\[3\sqrt{108} -2\sqrt{18} -3\sqrt{48}\]

OpenStudy (anonymous):

i need help solving it

OpenStudy (anonymous):

3*10.38-2*4.24-3*6.92

OpenStudy (anonymous):

or you can square the coefficients and multiply them my the numbers in the radicals and solve.

OpenStudy (anonymous):

i dont get it?

OpenStudy (anonymous):

This might help: 108 = 36*3, 18 = 9*2, 48 = 16*3. Can you see how to use that?

OpenStudy (anonymous):

\[6(\sqrt{3}-\sqrt{2})\]